# DFPE: A Diverse Fingerprint Ensemble for Enhancing LLM Performance

### v1

@@ 1 Introduction
Ensembling multiple LLMs provides a promising avenue to overcome these limitations Tekin et al. (2024); Jiang et al. (2023b); Mavromatis et al. (2024); Xu et al. (2024b). By leveraging the complementary strengths of different models, an ensemble can achieve greater accuracy, robustness, and adaptability than any single model alone Lu et al. (2024). However, maximizing the potential of an ensemble requires careful selection and integration of its models. Key challenges include identifying models that offer complementary strengths, preserving a diverse set of solution strategies, adapting to subject-specific difficulties, and efficiently aggregating predictions.
@@ 2 Related Work
A prominent direction involves assembling multiple models to capitalize on their unique abilities, as seen in LLM-TOPLA Tekin et al. (2024), PackLLM Mavromatis et al. (2024), and SweetSpan Xu et al. (2024b), each of which coordinates output information at various granularities (e.g., token- or span-level). While these techniques often produce strong results, they may also introduce extra training or inference overhead. Similar concepts appear in ZOOTER Lu et al. (2023), which dynamically routes queries to suitable models, and LoRA Ensembles Wang et al. (2023), which uses parameter-efficient fine-tuning to merge model capabilities. However, approaches that rely heavily on additional training or fine-tuning can be less practical when resources are limited or when rapid experimentation is required.
Other methods seek to remain fully or mostly training-free. Boosted Prompts Pitis et al. (2023) and CAPE Jiang et al. (2023c), for instance, forgo further training in favor of leveraging pretrained models and calibrating their outputs. This philosophy aligns with EVA Xu et al. (2024a) and DeePEn Huang et al. (2024) in that they preserve model diversity or transform model outputs at an inference or validation stage without extensive fine-tuning. However, many of these solutions do not adapt well to subject-wise performance variations, which becomes crucial for wide-ranging benchmarks like MMLU.
@@ 3.2 Formal Description
For each model $M_{i}$ and subject $S_{k}$, we produce a fingerprint vector $\mathbf{f}_{i,k}$ summarizing the model’s response behavior on validation data. For example, we can embed each output (e.g., using a pre-trained sentence embedding model) and average the embeddings to create the fingerprint. Once we obtain $\{\mathbf{f}_{i,k}\}_{i=1}^{N}$, we apply DBSCAN clustering with cosine similarity to group models into clusters based on their response patterns.
@@ 4.1 Dataset
We evaluate our method on the MMLU benchmark Hendrycks et al. (2020), which consists of 57 subjects spanning a diverse range of disciplines, including STEM fields, humanities, and social sciences. Each subject comprises multiple-choice questions that vary in difficulty. The given benchmark data were partitioned into a few-shot validation set for model selection and fingerprint generation, and a separate test set for the final evaluation.
@@ 4.2 Base Models
Our ensemble is constructed from a pool of LLMs differing in architecture, size, and training methodology, including GLM-4-9B-chat GLM et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024), Qwen2.5-3B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), Phi-3.5-mini-instruct Abdin et al. (2024), Llama-3.1-8B-Instruct Dubey et al. (2024), Gemma-2-9B Team (2024), Apollo2-7B Zheng et al. (2024), Starling-LM-7B-alpha Zhu et al. (2023), and Yi-1.5-6B Young et al. (2024). These models are detailed in Table 2. The best single model in our experiments achieves 70.76% overall-accuracy on MMLU, with a discipline-accuracy of 69%. Other models range from approximately 61.5% to 70.5% overall-accuracy, providing a rich diversity of performance levels. This variability creates an opportunity to exploit complementary strengths: by carefully selecting and weighting models via quantile-based filtering and clustering, we aim to surpass the capabilities of any single model.
@@ 4.3 Implementation Details
Few-Shot Validation for Selection and Weighting:
For each subject, given validation questions (approximately 10% of the test set) to measure model accuracy. These validation accuracies guide the quantile threshold, clustering-based selection, and the adaptive weighting scheme. By utilizing this few-shots validation approach rather than additional training or fine-tuning, our method maintains computational efficiency while being compatible with any pre-trained LLM, whether accessed through local deployment or API endpoints.
Clustering and Quantile Thresholding:
For each subject, we embed each model’s validation responses to form a fingerprint vector. We then cluster these fingerprints using DBSCAN with cosine similarity. The DBSCAN epsilon parameter is chosen based on empirical tuning to ensure stable, meaningful clusters. We select a quantile parameter $q$ = 0.05 based on initial validation performance. Models meeting or exceeding the $q$-quantile accuracy threshold are retained. This filtering guarantees that only models demonstrating a baseline level of competence for that subject remain candidates for the ensemble.
Adaptive Weighting and Hyperparameters:
We apply exponential scaling to each retained model’s per-subject validation accuracy. Parameters such as the AccuracyFactor and quantile thresholds are determined via validation-based exploration. For example, setting the AccuracyFactor ($\gamma$) to 5.0 and choosing $q=0.05$ consistently yields strong performance. After weighting models according to their accuracies, we aggregate their predictions with a weighted voting mechanism, producing final ensemble predictions without extra fine-tuning or complex training procedures.
@@ 5.5 Discussion
Our extensive experimental results provide several key insights into the effectiveness and practical implications of DFPE. We organize our discussion around three main aspects: performance characteristics, trade-offs, and practical considerations.
@@
Quantile Threshold: A low threshold (0.05) maximizes accuracy by including more models but increases computational cost. Higher thresholds (0.15-0.20) offer a more balanced trade-off, reducing the ensemble size while maintaining most performance benefits.
AccuracyFactor: Higher values (around 5.0) effectively emphasize stronger models’ contributions, but excessive values risk over-reliance on top performers, potentially reducing the benefits of ensemble diversity.
DBSCAN Epsilon: Smaller values produce more fine-grained clusters, preserving subtle model differences, while larger values lead to more aggressive grouping and smaller ensembles.
@@
Balanced Mode: Employs higher quantile thresholds and larger clustering epsilon, reducing ensemble size to 5-7 models while maintaining performance within 0.5-1.0
Efficient Mode: Further reduces model count through more aggressive filtering, suitable for resource-constrained environments while still outperforming single models.
@@
Limitations
While DFPE demonstrates significant improvements, several limitations warrant discussion:
Cold-start Performance: The method requires validation data for initial model selection and weighting, which may not always be available for new subjects or domains.
Computational Overhead: In its highest-accuracy configuration, the method requires running multiple models, which may be impractical for some real-time applications.
Model Diversity Requirement: The effectiveness of DFPE depends on having access to a diverse pool of models with complementary strengths.

### v2

@@ 1 Introduction
Ensembling multiple LLMs provides a promising avenue to overcome these limitations Tekin et al. (2024); Jiang et al. (2023b); Mavromatis et al. (2024); Xu et al. (2025). By leveraging the complementary strengths of different models, an ensemble can achieve greater accuracy, robustness, and adaptability than any single model alone Lu et al. (2024). However, maximizing the potential of an ensemble requires careful selection and integration of its models. Key challenges include identifying models that offer complementary strengths, preserving a diverse set of solution strategies, adapting to subject-specific difficulties, and efficiently aggregating predictions.
@@ 2 Related Work
A prominent direction involves assembling multiple models to capitalize on their unique abilities, as seen in LLM-TOPLA Tekin et al. (2024), PackLLM Mavromatis et al. (2024), and SweetSpan Xu et al. (2025), each of which coordinates output information at various granularities (e.g., token- or span-level). While these techniques often produce strong results, they may also introduce extra training or inference overhead. Similar concepts appear in ZOOTER Lu et al. (2023), which dynamically routes queries to suitable models, and LoRA Ensembles Wang et al. (2023), which uses parameter-efficient fine-tuning to merge model capabilities. However, approaches that rely heavily on additional training or fine-tuning can be less practical when resources are limited or when rapid experimentation is required.
Other methods seek to remain fully or mostly training-free. Boosted Prompts Pitis et al. (2023) and CAPE Jiang et al. (2023c), for instance, forgo further training in favor of leveraging pretrained models and calibrating their outputs. This philosophy aligns with EVA Xu et al. (2024) and DeePEn Huang et al. (2024) in that they preserve model diversity or transform model outputs at an inference or validation stage without extensive fine-tuning. However, many of these solutions do not adapt well to subject-wise performance variations, which becomes crucial for wide-ranging benchmarks like MMLU.
@@ 3.2 Formal Description
For each model $M_{i}$ and subject $S_{k}$, we produce a fingerprint vector $\mathbf{f}_{i,k}$ summarizing the model’s response behavior on validation data. We use the pre-trained sentence embedding model all-MiniLM-L6-v2 Face (2023) to embed each output and average the embeddings to create the fingerprint.. Once we obtain $\{\mathbf{f}_{i,k}\}_{i=1}^{N}$, we apply DBSCAN Ester et al. (1996) clustering with cosine similarity to group models into clusters based on their response patterns.
@@ 4.1 Dataset
We evaluate our method on the MMLU benchmark Hendrycks et al. (2020), which consists of 57 subjects spanning a diverse range of disciplines, including STEM fields, humanities, and social sciences. Each subject comprises multiple-choice questions in English, which vary in difficulty. The given benchmark data is partitioned into a few-shot validation set for model selection and fingerprint generation, and a separate test set for the final evaluation. Overall, there are 14,079 test samples and 1,540 validation samples. The MMLU dataset is distributed under the MIT License, which allows for free use, modification, and distribution as long as the original copyright notice and license terms are maintained.
@@ 4.2 Base Models
Our ensemble is constructed from a pool of up to 9B parameters open LLMs differing in architecture, size, and training methodology, including GLM-4-9B-chat GLM et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024), Qwen2.5-3B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), Phi-3.5-mini-instruct Abdin et al. (2024), Llama-3.1-8B-Instruct Dubey et al. (2024), Gemma-2-9B Team et al. (2024), Apollo2-7B Zheng et al. (2024), Starling-LM-7B-alpha Zhu et al. (2024), and Yi-1.5-6B Young et al. (2024). These models are detailed in Table 2. The best single model in our experiments achieves 70.76% overall-accuracy on MMLU, with a discipline-accuracy of 69%. Other models range from approximately 61.5% to 70.5% overall-accuracy, providing a rich diversity of performance levels. This variability creates an opportunity to exploit complementary strengths: by carefully selecting and weighting models via quantile-based filtering and clustering, we aim to surpass the capabilities of any single model.
@@ 4.3 Implementation Details
Clustering and Quantile Thresholding:
For each subject, we embed each model’s validation responses to form a fingerprint vector. We then cluster these fingerprints using DBSCAN with cosine similarity. The DBSCAN epsilon parameter is chosen based on empirical tuning to ensure stable, meaningful clusters. We select a quantile parameter $q$ = 0.05 based on initial validation performance. Models meeting or exceeding the $q$-quantile accuracy threshold are retained.
Adaptive Weighting and Hyperparameters:
We apply exponential scaling to each retained model’s per-subject validation accuracy. Parameters such as the AccuracyFactor and quantile thresholds are determined via validation-based exploration. For example, setting the AccuracyFactor ($\gamma$) to 5.0 and choosing $q=0.05$ consistently yields strong performance.
Few-Shot Validation for Selection and Weighting:
For each subject, given validation questions (approximately 10% of the test set) to measure model accuracy. These validation accuracies guide the quantile threshold, clustering-based selection, and the adaptive weighting scheme.
Hardware Details:
All experiments were conducted on RTX 6000 GPU, which provided sufficient computational resources for up to 9B parameters LLMs.
@@ 5.5 Discussion
Our extensive experimental results provide several key insights into the effectiveness and practical implications of DFPE. We organize our discussion around three main aspects: performance characteristics, Model Pool Homogeneity, parameter insensitivity, and trade-offs for practical considerations.
@@
Model Pool Homogeneity
It is important to note that the method’s effectiveness hinges on the diversity present within the pool of base LLMs. Designed to capitalize on complementary strengths and varied perspectives from heterogeneous models, its benefits could diminish with a more homogeneous pool – for instance, models with similar architectures, training data, or biases. In such cases, fingerprint clustering may become less informative, and the ensemble would mainly function as averaging largely redundant model outputs. Consequently, performance gains achieved over a strong single model might be less pronounced, as the ensemble would lack the advantage of integrating genuinely diverse and contrasting insights. Future investigations could explore the method’s performance under varying degrees of model pool homogeneity to further delineate its applicability and robustness.
@@
Quantile Threshold: Low thresholds (e.g., 0.05) maximize accuracy by including more models, increasing cost. Higher thresholds (0.15-0.20) offer a balanced approach, reducing ensemble size while maintaining performance.
AccuracyFactor: Higher values (around 5.0) prioritize strong models, but can risk over-reliance and reduce the benefits of ensemble diversity.
DBSCAN Epsilon: Small values create finer clusters, preserving subtle model differences. Large values lead to aggressive grouping and smaller ensembles.
@@
Balanced Mode: Employs higher quantile thresholds and larger clustering epsilon, reducing ensemble size to 5-7 models while maintaining performance within 0.5-1.0% of optimal.
Efficient Mode: Further reduces model count through more aggressive filtering, suitable for resource-constrained environments.
@@ 6 Conclusion
Limitations
While DFPE demonstrates significant improvements, several limitations warrant discussion: Dependence on Few-Shot Validation Data DFPE relies on few-shot validation data for initial model selection and weighting. Computational Overhead: Running multiple models requires more computational resources than using a single small model. Model Diversity Requirement: The approach benefits from having a diverse pool of models with complementary strengths, and its effectiveness may reduced with homogeneous model sets; Task-Specificity: The current implementation is confined to multiple-choice question answering.
Acknowledgements
We used ChatGPT-4o for editing the language and refining the presentation of the text in this paper. The authors affirm that all research content and ideas are their own, and they take full responsibility for the final submitted manuscript.